OpenStudy (anonymous):

Problem:5 over 10%x55=____ Choices:A. .00275 B. .0275 C. .275 D. 2.75

OpenStudy (anonymous):

(5/10)% = 1/2% = 1/2 * 1% 1 % of 55 is .55, half of that is .275

OpenStudy (anonymous):

i dont get it. which part is the answer ??? lol